# PromptEcho: Annotation-Free Reward from Vision-Language Models for Text-to-Image Reinforcement Learning

### 14 Apr 2026 v1

@@ Abstract
Reinforcement learning (RL) can improve the prompt following capability of text-to-image (T2I) models, yet obtaining high-quality reward signals remains challenging: CLIP Score is too coarse-grained, while VLM-based reward models (e.g., RewardDance) require costly human-annotated preference data and additional fine-tuning. We propose PromptEcho, a reward construction method that requires no annotation and no reward model training. Given a generated image and a guiding query, PromptEcho computes the token-level cross-entropy loss of a frozen VLM with the original prompt as the label, directly extracting the image-text alignment knowledge encoded during VLM pretraining. The reward is deterministic, computationally efficient, and improves automatically as stronger open-source VLMs become available. For evaluation, we develop DenseAlignBench, a benchmark of concept-rich dense captions for rigorously testing prompt following capability. Experimental results on two state-of-the-art T2I models (Z-Image and QwenImage-2512) demonstrate that PromptEcho achieves substantial improvements on DenseAlignBench (+26.8pp / +16.2pp net win rate), along with consistent gains on GenEval, DPG-Bench, and TIIFBench without any task-specific training. Ablation studies confirm that PromptEcho comprehensively outperforms inference-based scoring with the same VLM, and that reward quality scales with VLM size. We will open-source the trained models and the DenseAlignBench.
@@ 3.1 PromptEcho Reward
If image $x$ faithfully depicts prompt $c$ (e.g., correctly rendering all objects, attributes, and spatial relationships), the VLM should “recite” each token of $c$ with high probability after observing the image—the image echoes the prompt—yielding a high reward. Conversely, omissions or semantic deviations in the image lower the token-level likelihood and thus the reward. This formulation works because the reward is exactly the VLM’s pretraining objective (token-level cross-entropy), directly tapping into the image-text alignment knowledge acquired from massive data—without any additional conversion layers, regression heads, or fine-tuning. We analyze this “understanding vs. reasoning” distinction further in Section 5.1.
@@ 4.1 Experimental Setup
DenseAlignBench (ours). To rigorously test the prompt following capability of state-of-the-art T2I models on concept-rich, detailed descriptions, we construct DenseAlignBench—a test set of dense captions with rich concepts, relationships, and fine-grained details. DenseAlignBench consists of 2000 test captions separated from our internal dataset; these captions are not in the training set but share the same source and distribution as the training data. We use Gemini-3-flash-preview (Google DeepMind, 2025) for pairwise evaluation on the prompt following dimension, directly reflecting the method’s effectiveness. We will open-source the DenseAlignBench.
@@ 4.3.1 PromptEcho vs InferScore
To directly compare PromptEcho (extracting VLM “understanding”) with InferScore (leveraging VLM “reasoning”), we design a controlled ablation: both use the identical VLM (Qwen3-VL-32B) and the same base model (Z-Image), differing only in reward construction. Results are shown in Tables 56.
@@
The results reveal a stark contrast. On DenseAlignBench, PromptEcho achieves +26.8pp net advantage while InferScore falls below the baseline (-3.3pp). On public benchmarks, PromptEcho consistently outperforms the baseline (GenEval +6.5pp, DPG-Bench +1.02, TIIFBench +3.6/+5.8pp), while InferScore shows negligible improvement or degradation (GenEval +1pp, DPG-Bench -0.04, TIIFBench short -2.5pp). This confirms that extracting VLM “understanding” via the loss function is far more effective than leveraging VLM “reasoning” via inference (see Section 5.1).
@@ 5 Discussion
5.1 Understanding vs. Reasoning
InferScore relies on the VLM’s reasoning ability—explicitly judging and scoring image-text alignment through autoregressive generation, a process limited by inference-time hallucinations and sampling randomness. PromptEcho instead extracts the VLM’s understanding ability: the implicit image-text alignment knowledge encoded in the pretrained weights, accessed through the same cross-entropy loss used during pretraining. Because next-token prediction accuracy is typically more robust than explicit reasoning chains, PromptEcho provides more stable and reliable reward signals—as confirmed by the large performance gap in our ablation (Section 4.3.1).

### 22 Apr 2026 v2

@@ Abstract
Reinforcement learning (RL) can improve the prompt following capability of text-to-image (T2I) models, yet obtaining high-quality reward signals remains challenging: CLIP Score is too coarse-grained, while VLM-based reward models (e.g., RewardDance) require costly human-annotated preference data and additional fine-tuning. We propose PromptEcho, a reward construction method that requires no annotation and no reward model training. Given a generated image and a guiding query, PromptEcho computes the token-level cross-entropy loss of a frozen VLM with the original prompt as the label, directly extracting the image-text alignment knowledge encoded during VLM pretraining. The reward is deterministic, computationally efficient, and improves automatically as stronger open-source VLMs become available. For evaluation, we develop DenseAlignBench, a benchmark of concept-rich dense captions for rigorously testing prompt following capability. Experimental results on two state-of-the-art T2I models (Z-Image and QwenImage-2512) demonstrate that PromptEcho achieves substantial improvements on DenseAlignBench (+26.8pp / +16.2pp net win rate), along with consistent gains on GenEval, DPG-Bench, and TIIFBench without any task-specific training. Ablation studies confirm that PromptEcho comprehensively outperforms inference-based scoring with the same VLM, and that reward quality scales with VLM size. We have open-sourced the trained models and the DenseAlignBench (https://github.com/roooobotx/prompt_echo).
@@ 3.1 PromptEcho Reward
If image $x$ faithfully depicts prompt $c$ (e.g., correctly rendering all objects, attributes, and spatial relationships), the VLM should “recite” each token of $c$ with high probability after observing the image—the image echoes the prompt—yielding a high reward. Conversely, omissions or semantic deviations in the image lower the token-level likelihood and thus the reward. This formulation works because the reward is exactly the VLM’s pretraining objective (token-level cross-entropy), directly tapping into the image-text alignment knowledge acquired from massive data—without any additional conversion layers, regression heads, or fine-tuning. We analyze this mechanism further in Section 5.1.
@@ 4.1 Experimental Setup
DenseAlignBench (ours). To rigorously test the prompt following capability of state-of-the-art T2I models on concept-rich, detailed descriptions, we construct DenseAlignBench—a test set of dense captions with rich concepts, relationships, and fine-grained details. DenseAlignBench consists of 2000 test captions separated from our internal dataset; these captions are not in the training set but share the same source and distribution as the training data. We use Gemini-3-flash-preview (Google DeepMind, 2025) for pairwise evaluation on the prompt following dimension, directly reflecting the method’s effectiveness. We have open-sourced the DenseAlignBench [^1] https://github.com/roooobotx/prompt_echo.
@@ 4.3.1 PromptEcho vs InferScore
To directly compare PromptEcho (computing continuous log-likelihood via the pretraining loss) with InferScore (generating discrete scores via autoregressive inference), we design a controlled ablation: both use the identical VLM (Qwen3-VL-32B) and the same base model (Z-Image), differing only in reward construction. Results are shown in Tables 56.
@@
The results reveal a stark contrast. On DenseAlignBench, PromptEcho achieves +26.8pp net advantage while InferScore falls below the baseline (-3.3pp). On public benchmarks, PromptEcho consistently outperforms the baseline (GenEval +6.5pp, DPG-Bench +1.02, TIIFBench +3.6/+5.8pp), while InferScore shows negligible improvement or degradation (GenEval +1pp, DPG-Bench -0.04, TIIFBench short -2.5pp). The core reason for InferScore’s failure is that its discrete scoring mechanism prevents the VLM from distinguishing subtle differences in prompt following across multiple images generated from the same prompt with different seeds—frequently assigning identical scores to all images, effectively nullifying the reward signal. In contrast, PromptEcho outputs continuous log-likelihood values that naturally capture fine-grained distinctions (see Section 5.1).
@@ 5 Discussion
5.1 Pretraining Loss vs. Autoregressive Scoring
InferScore lets the VLM explicitly score images through autoregressive generation, a process subject to hallucinations and sampling randomness. More critically, its discrete scoring mechanism prevents the VLM from distinguishing subtle differences in prompt following across multiple images generated from the same prompt with different seeds—frequently assigning identical scores to all images, effectively nullifying the reward signal. PromptEcho instead directly computes continuous log-likelihood values via the VLM’s pretraining cross-entropy loss, requiring no autoregressive generation. The result is fully deterministic and naturally captures fine-grained distinctions—as confirmed by the large performance gap in our ablation (Section 4.3.1).